You have arranged that the magnetic field in a particular region of space is due North with a value of 0.0100 T. An electron enters the field traveling to the West with a speed of 6.0 percent of the speed of light. As a result, the electron experiences a magnetic force which is upwards.

(a) What is the strength of this magnetic force? Give your answer in femtonewtons. (NOTE: The prefix "femto" represents 10^(-15).) fN

(b) What is the amount of the resulting acceleration of the electron? Get the mass of an electron from the inside back cover of your textbook. Give your answer in Pm/s2 (NOTE: the prefix "P" is for "peta", which represents 10^(15).) Pm/s2

(c) What will be the result of the acceleration that you have calculated in part (b)?

a. Both the speed and direction of the moving electron will change.

b.The direction of the moving electron will change but its speed will remain constant.

c.The speed of the moving electron will change but its direction of motion will remain constant..
Answer #1

(a)

F=evB=(1.6\times 10^{-19}\times 0.06\times 3\times 10^8 \times 0.0100)\: N = 28.8\: fN

(b)

a=\frac {F}{m}=\frac {28.8\times 10^{-15}}{9.1\times 10^{-31}}\: m/s^2 =31.648\: pm/s^2

(c)

The direction of the moving electron will change but its speed will remain constant.